# Embedded coherent structures from MHD to sub-ion scales in turbulent solar wind at 0.17 AU

### v2

@@ Abstract
We study intermittent coherent structures in solar wind magnetic turbulence from MHD to kinetic plasma scales using Parker Solar Probe data during its first perihelion (at 0.17 au), when the satellite was in the Alfvénic slow wind of $\sim 340$ km/s. The coherent structures are energetic events localized in time and covering wide range of scales. We detect them using Morlet wavelets. For the first time, we apply a multi-scale analyses in physical space to study these structures. At MHD scales within the inertial range, times scales $\tau\in(1,10^{2})$ s, we find (i) current sheets including switchback boundaries and (ii) Alfvén vortices. Within these events, there are embedded structures at smaller scales: typically Alfvén vortices at ion scales, $\tau\in(0.08,1)$ s, and a compressible vortices at sub-ion scales, $\tau\in 8(10^{-3},10^{-2})$ s. The number of coherent structures grows toward smaller scales: we observe $\sim 200$ events during 5 h time interval at MHD scales, $\sim 10^{3}$ events ai ion scales and $\sim 10^{4}$ events at sub-ion scales. In general, there are multiple structures of ion and sub-ion scales embedded within one MHD structure. There are also examples of ion and sub-ion scales structures outside of MHD structures.
To quantify the relative importance of different type of structures, we do a statistical comparison of the observed structures with the expectations of models of the current sheets and vortices. This comparison is based on amplitude anisotropy of magnetic fluctuations within the structures. The results show the dominance of Alfvén vortices at all scales in contrast to the widespread view of dominance of current sheets. This means that Alfvén vortices are important building blocs of solar wind turbulence.
@@ 1 Introduction
Solar wind fluctuations cover a broad range of scales: from macroscopic scales, where the energy is injected into the MHD turbulent cascade, to micro-scales, where kinetic effects play important role, and the energy is dissipated. The dissipation mechanism has not been understood yet. Numerical simulations indicate that dissipation occurs inhomogeneously (Wan et al., 2012; Karimabadi et al., 2013; Zhdankin et al., 2013; Kuzzay et al., 2019). Regions of increased heating in the solar wind correlate with observations of coherent structures (Osman et al., 2011; Wu et al., 2013; Chasapis et al., 2015; Sioulas et al., 2022). Coherent structures can be defined as high-amplitude, stable, localized in space events with phase coherence over its spatial extent (Hussain, 1986; Fiedler, 1988; Veltri, 1999; Bruno et al., 2001; Mangeney, 2001; Farge & Schneider, 2015; Alexandrova, 2020).
@@
Different types of coherent structures are observed in the solar wind at different scales. Large scale flux tubes and flux ropes cover energy containing scales and the inertial range (e.g. Moldwin et al., 2000; Feng et al., 2008; Borovsky, 2008; Janvier et al., 2014; Zhao et al., 2020). Current sheets are usually observed at small scales of the inertial range and at ion scales (e.g. Siscoe et al., 1968; Burlaga, 1969; Salem, 2000; Knetter et al., 2004; Tsurutani et al., 2011; Lion et al., 2016; Perrone et al., 2016; Artemyev et al., 2019). Recent Solar Orbiter observations reveal embedded ion scale flux rope in a bifurcated current sheet (Eastwood et al., 2021). Alfvén vortices have been identified at MHD scales and at ion scales (Verkhoglyadova et al., 2003; Roberts et al., 2016; Lion et al., 2016; Perrone et al., 2016, 2017). Compressible structures, such as magnetic holes (e.g. Turner et al., 1977; Stevens & Kasper, 2007; Volwerk et al., 2020)), solitons and shocks (Perrone et al., 2016; Salem, 2000) are observed at the end of the inertial range and at ion scales.
@@
Coherent structures contribute significantly to the magnetic turbulent spectrum in the solar wind. Li et al. (2011) show that in the presence of current sheets, the inertial range spectrum is closer to the Kolmogorov scaling, $-5/3$, while without current sheets, the spectrum is closer to the Iroshnikov-Kraichnan scaling, $-3/2$. In a case study of a fast wind stream by Lion et al. (2016), the contribution of coherent structures to magnetic field spectrum is up to 40 % from inertial range down to ion scales. Therefore, coherent structures are energetically important elements of solar wind turbulence.
@@
Much less observations of coherent structures exist at sub-ion scales. Cluster/STAFF allows to measure sub-ion scale fluctuations at 1 au. Greco et al. (2016) studied ion scales current sheets and showed presence of a number of smaller ones at sub-ion scales. This has been done using the partial variance of increments (PVI) method (Greco et al., 2018), which is appropriate to detect planar structures with important field rotation (e.g., see the discussion in Lion et al., 2016). Another study with Cluster, but applying Morlet wavelets show embedded Alfvén vortex type fluctuations at sub-ion scales in a current sheet at ion scales (Jovanović et al., 2018). An analytic model of a chain of Alfvén vortices embedded in the current sheet is developed by authors.
@@
First multi-satellite observation of these cylindrical structures have been done in the Earth’s magnetosheath with the Cluster mission (Alexandrova et al., 2006; Alexandrova, 2008). Cassini measurements indicate the presence of such structures in the Kronian magnetosheath as well (Alexandrova & Saur, 2008). Signatures of Alfvén vortices in the solar wind using one satellite have been shown by Verkhoglyadova et al. (2003) and Lion et al. (2016). Roberts et al. (2016) and Perrone et al. (2016, 2017) confirmed the existence of Alfven vortices in the solar wind with 4 satellites of Cluster. Wang et al. (2019) investigated the kinetic effects within an Alfvén vortex thanks to MMS measurements in the Earth’s magnetosheath.
@@
In the present paper, we study magnetic turbulent fluctuations from MHD inertial range to sub-ion scales with Parker Solar Probe (PSP) data at 0.17 au. Using the Morlet wavelet transform, which is a good compromise between time and frequency resolution, we detect intermittent events which cover a wide range of scales. We show that these events correspond to embedded multi scale structures, from MHD to sub-ion scales. Then, we study in more details nature of these structures, which cover the whole turbulent cascade.
@@
The article is organized as follows: Section 2 describes the PSP data used in the analysis. In Section 3 magnetic spectral properties are discussed. Section 4 is dedicated to the detection method of coherent structures. Section 5 presents several theoretical models of the structures we think we cross by PSP at different scales. Then we study the sensitivity of minimum variance results for different spacecraft trajectories across the model structures for different noise levels. In section 6 we describe few examples of detected structures at MHD, ion and sub-ion scales simultaneously. Section 7 describes statistical study of the observed coherent events during 5h of the 1st PSP encounter at MHD, ion and sub-ion scales. In section 8 we summarize and discuss the results.
@@ 2 Data
Figure 1: Overview of the solar wind data during the first perihelion of PSP at 0.17 au on November 6, 2018, between 00:00 and 05:00 UT. From top to bottom: (a) the magnetic field components in RTN reference frame and the magnetic field modulus, (b) proton bulk velocity components, (c) angle between the magnetic field and the proton bulk velocity, (d) electron density, (e) proton temperature (in blue) and core electron temperatures (in black), (f) core electron and proton plasma beta, (g) ion characteristic scales, $d_{i}$ and $\rho_{i}$.
We analyze 5 hours time interval during the first perihelion, on November 6, 2018, [00:00, 05:00] UT, when the spacecraft at the distance of 0.17 au from the Sun measured the solar wind emerging from the small equatorial coronal hole (Bale et al., 2019; Kasper et al., 2019). The magnetic field during the chosen time interval is particularly highly-disturbed due to the presence of high-amplitude structures (including switchbacks, Bale et al., 2019; Perrone et al., 2020). The duration of the chosen interval is long enough to resolve the inertial range of MHD turbulence, but not too long, so that the PSP is magnetically connected to the same coronal hole and the PSP position is nearly at the same radial distance from the Sun.
We use the merged magnetic field measurements of two magnetometers: FIELDS/Fluxgate Magnetometer and Search Coil (Bale et al., 2016; Bowen et al., 2020). These data have $3.4$ ms time resolution, which allows us to resolve a wide range of scales, going from MHD inertial range to sub-ion range. Due to the Search Coil anomaly at one of its axis (happened on March 2019), the full merged vector of the magnetic field is accessible only for the first perihelion (Bowen et al., 2020). Figure 1(a) shows the magnetic field magnitude $B(t)$ in black and three components in RTN coordinate frame in color.
To characterize ion plasma parameters we use SWEAP/SPC Faraday cup instrument (Kasper et al., 2016). Proton velocity $\mathbf{V}$, estimated from the 1st moment of the distribution function, is shown at Figure 1(b). The mean proton velocity is nearly radial $\langle\mathbf{V}\rangle=(330,70,8)$ km/s. The mean angle with the radial direction is $\langle\Theta_{\mathbf{V},\mathbf{R}}\rangle=14^{\circ}$. The magnetic field vector fluctuates around $\langle\mathbf{B}\rangle=(-47,63,-5)$ nT (Figure 1(a)). It’s magnitude is nearly constant $|\mathbf{B}|=98\pm 5$ nT. The angle between the magnetic field and velocity changes from $20^{\circ}$ to $180^{\circ}$ as shown in Figure 1(c), with a dominance around orthogonal crossings of the magnetic field, $\langle\Theta_{\mathbf{B},\mathbf{V}}\rangle=107^{\circ}$ with standard deviation $\sigma(\Theta_{\mathbf{B},\mathbf{V}})=26^{\circ}$.
@@
Proton temperature $T_{p}$ is estimated from the second moment of the distribution function measured by SWEAP/SPC instrument. QTN electron core temperature $T_{e,core}$ and proton temperature $T_{p}$ are shown in Figure 1(e).
We consider solar wind fluctuations over a wide range of timescales. The minimum time scale is determined by the temporal resolution of the magnetic data $dt=3.4\cdot 10^{-3}$ s, and the maximum by the total duration of the time interval $T=5$ h. The corresponding range of the spatial scales estimated using Taylor hypothesis with $V=340$ km/s, $\ell\in[1,10^{5}]$ km are much larger than the Debye length $\lambda_{D}\simeq 2$ m, then the plasma is quasi-neutral. During the analysed time period, alpha particle abundance $A_{\alpha}=N_{\alpha}/N_{p}<5\%$ is negligible (Kasper et al., 2007; Alterman & Kasper, 2019). The quasi-thermal noise spectroscopy provides more accurate measurement of the density than particle detectors, so we use $N_{p}=N_{e}$ and calculate proton plasma beta $\beta_{p}$ using the electron density: $\beta_{p}=N_{e}kT_{p}/(B^{2}/2\mu_{0})$, with $\mu_{0}$ being the magnetic permeability. Plasma beta for core electrons is defined as $\beta_{e,core}=N_{e}kT_{e,core}/(B^{2}/2\mu_{0})$. Both plasma $\beta$ parameters are well below unity as shown in Figure 1(f). Figure 1 (g) shows time variation of the ion characteristic spatial scales $d_{i}=c/\omega_{pi}$, where $\omega_{pi}=\sqrt{4\pi N_{p}e^{2}/m_{p}}$ and $\rho_{i}=V_{th,\perp}/\omega_{ci}$, where $V_{th,\perp}$ is the perpendicular proton thermal speed, $\omega_{ci}=eB/m_{p}c$ is the proton gyrofrequency.
3 MHD, ion and sub-ion ranges identification
Figure 2: From top to bottom: magnetic field total spectrum $S_{total}$ in orange and magnetic field modulus spectrum computed with the parallel fluctuations spectrum $S_{\parallel}$ (Equation (5)) (a), and the ratio $S_{\parallel}/S_{total}$ (b). The vertical lines show the characteristic ion scales: ion cyclotron frequency $f_{ci}$ (in blue), and the frequencies computed with the Doppler shifted ion gyroradius $f_{\rho i}$ (in black) and the Doppler shifted ion inertial length $f_{di}$ (in grey). The frequency ranges are highlighted: MHD in red, ion scales in green and sub-ion scales in blue.
First we describe the magnetic field spectral properties of the analysed time interval. We apply wavelet transform with Morlet mother function (Torrence & Compo, 1998):
@@ 3 MHD, ion and sub-ion ranges identification
where $\omega_{0}=6$ is the angular frequency of oscillations in the mother function (with normalized time). The wavelet transform of the magnetic field component $B_{i}(t)$ is defined as the convolution of $B_{i}(t)$ with scaled, translated and normalized $\psi_{0}(t)$ to have mother function $\psi$ with unit energy:
@@
where the sign ${}^{\ast}$ indicates complex conjugate.
Wavelet coefficients are influenced by the edge effects. Cone of influence (COI) curve separates the region of scales where edge effects become important as a function of time. To avoid this edge effect we consider a maximum scale equal to $\tau_{max}=10^{3}$ s. The intercept of $\tau_{max}$ with COI curve determines the time sub-interval $T^{\prime}=$ [00:22:49, 04:37:11] UT, where wavelet coefficients at the scales $\tau<\tau_{max}$ are not influenced by the edge effect.
@@
where $\delta t=0.008$ s is the time-step of the PSP merged magnetic field data. The relation between Fourier frequencies $f$ and time scales $\tau$ is $f\simeq 1/\tau$ for the Morlet wavelets with $\omega_{0}=6$. Figure 2(a), blue line, shows the PSD of compressive magnetic fluctuations. Compressive fluctuations are approximated here by the variation of magnetic field modulus. Indeed, this approximation is valid if the level of the fluctuations is significantly lower than the mean field $B_{0}$, i.e., $\delta B/B_{0}\ll 1$ (Perrone et al., 2016):
@@
In the inertial range and at higher frequencies the condition $\delta B/B_{0}\ll 1$ is valid. So we calculate the parallel PSD, $S_{\parallel}(\tau)$, as it was done in Perrone et al. (2016):
@@
The ratio of compressible fluctuations to the total power spectral density $S_{\parallel}/S_{total}$ is shown in Figure 2(b). In the inertial range, parallel magnetic fluctuations are much less energetic than perpendicular ones ($\delta B_{\|}\ll\delta B_{\perp}$), as is usually observed in the solar wind. At the sub-ion scales, the fraction of the parallel $S_{\parallel}(\tau)/S_{total}(\tau)$ increases up to $\simeq 0.2$, which is consistent with the results of Salem et al. (2012) at 1 au. The authors suggested that the observed spectral ratio can be explained by the presence of the kinetic Alfvén wave (KAW) cascade with nearly perpendicular wavevectors ($k_{\perp}\gg k_{\|}$). However, analyzing Cluster measurements (Lacombe et al., 2017) and 2D hybrid numerical simulation (Matteini et al., 2020) found that asymptotic compressibility value at sub-ion scales doesn’t match perfectly the KAW prediction. Finally, recent numerical simulations indicate that coherent structures, rather than waves, are energetically dominant on sub-ion scales (Papini et al., 2021).

### v4

@@ Abstract
We study intermittent coherent structures in solar wind turbulence from MHD to kinetic plasma scales using Parker Solar Probe data during its first perihelion (at 0.17 au) in the highly-Alfvénic slow solar wind. We detect coherent structures using Morlet wavelets. For the first time, we apply a multi-scale analysis in physical space. At MHD scales within the inertial range, times scales $\tau\in(1,10^{2})$ s, we find (i) current sheets including switchback boundaries and (ii) Alfvén vortices. Within these events are embedded structures at smaller scales: typically Alfvén vortices at ion scales, $\tau\in(0.08,1)$ s, and compressible vortices at sub-ion scales, $\tau\in 8(10^{-3},10^{-2})$ s. The number of coherent structures grows toward smaller scales: we observe $\sim 200$ events during a 5 h time interval at MHD scales, $\sim 10^{3}$ at ion scales, and $\sim 10^{4}$ at sub-ion scales. In general, there are multiple structures of ion and sub-ion scales embedded within one MHD structure. There are also examples of ion and sub-ion scale structures outside MHD structures.
To quantify the relative importance of different types of structures, we do a statistical comparison of the observed structures with the expectations of models of the current sheets and vortices. The results show the dominance of Alfvén vortices at all scales in contrast to the widespread view of the dominance of current sheets. This means that Alfvén vortices are important building blocks of Alfvénic solar wind turbulence.
@@ 1 Introduction
Solar wind fluctuations cover a broad range of scales: from macroscopic scales, where the energy is injected into the MHD turbulent cascade, to micro-scales, where kinetic effects play an important role, and the energy is dissipated. The dissipation mechanism has not been understood yet. Numerical simulations indicate that dissipation occurs inhomogeneously (Wan et al., 2012; Karimabadi et al., 2013; Zhdankin et al., 2013; Kuzzay et al., 2019). Regions of increased heating in the solar wind correlate with observations of coherent structures (Osman et al., 2011; Wu et al., 2013; Chasapis et al., 2015; Sioulas et al., 2022). Coherent structures can be defined as high-amplitude, stable, localized in space events with phase coherence over its spatial extent (Hussain, 1986; Fiedler, 1988; Veltri, 1999; Bruno et al., 2001; Mangeney, 2001; Farge & Schneider, 2015; Alexandrova, 2020).
@@
Different types of coherent structures are observed in the solar wind at different scales. Large-scale flux tubes and flux ropes cover energy-containing scales and the inertial range (e.g., Moldwin et al., 2000; Feng et al., 2008; Borovsky, 2008; Janvier et al., 2014; Zhao et al., 2020). Current sheets are usually observed at small scales of the inertial range and at ion scales (e.g., Siscoe et al., 1968; Burlaga, 1969; Salem, 2000; Knetter et al., 2004; Tsurutani et al., 2011; Lion et al., 2016; Perrone et al., 2016; Artemyev et al., 2019). Recent Solar Orbiter observations reveal embedded ion scale flux rope in a bifurcated current sheet (Eastwood et al., 2021). Alfvén vortices have been identified at MHD and ion scales (Verkhoglyadova et al., 2003; Roberts et al., 2016; Lion et al., 2016; Perrone et al., 2016, 2017). Recent numerical simulation shows that magnetic vortices emerge at the late stage of the reflection-driven turbulence (Meyrand et al., 2023). Compressible structures, such as magnetic holes (e.g., Turner et al., 1977; Stevens & Kasper, 2007; Volwerk et al., 2020)), solitons and shocks (Perrone et al., 2016; Salem, 2000) are observed at the end of the inertial range and at ion scales.
@@
Coherent structures contribute significantly to the magnetic turbulent spectrum in the solar wind. Li et al. (2011) show that in the presence of current sheets, the inertial range spectrum is closer to the Kolmogorov scaling, $-5/3$, while without current sheets, the spectrum is closer to the Iroshnikov-Kraichnan scaling, $-3/2$. In a case study of a fast wind stream by Lion et al. (2016), the contribution of coherent structures to the magnetic field spectrum is up to 40 % from the inertial range down to ion scales. Therefore, coherent structures are energetically important elements of solar wind turbulence.
@@
There are less observations of coherent structures on sub-ion scales. Cluster/STAFF allows to measure sub-ion scale fluctuations at 1 au. Greco et al. (2016) studied ion scale current sheets and showed the presence of many smaller ones at sub-ion scales. In this study the authors used the partial variance of increments (PVI) method (Greco et al., 2018), which is appropriate for detecting planar structures (e.g., see the discussion in Lion et al., 2016). Another study with Cluster, but applying Morlet wavelets, shows embedded Alfvén vortex-like fluctuations at sub-ion scales in a current sheet at ion scales (Jovanović et al., 2018). To explain this observation, the authors developed the analytic model of a chain of Alfvén vortices embedded in the current sheet. At electron scales there are also signatures of Alfvén-vortex like fluctuations detected by two Cluster satellites separated by 7 km only (Alexandrova, 2020; Alexandrova et al., 2020).
@@
The first multi-satellite observations of these cylindrical structures in space plasma were obtained with the Cluster mission in the Earth’s magnetosheath behind a quasi-perpendicular bow-shock (Alexandrova et al., 2006; Alexandrova, 2008). Cassini measurements also indicate the presence of such structures in the Kronian magnetosheath (Alexandrova & Saur, 2008). Signatures of Alfvén vortices in the solar wind using one satellite have been shown by Verkhoglyadova et al. (2003) and Lion et al. (2016). Roberts et al. (2016) and Perrone et al. (2016, 2017) confirmed the existence of Alfven vortices in the solar wind with 4 satellites of Cluster. Wang et al. (2019) investigated the kinetic effects within an Alfvén vortex thanks to MMS measurements in the Earth’s magnetosheath.
@@
In the present paper, we study magnetic turbulent fluctuations from MHD inertial range to sub-ion scales with Parker Solar Probe (PSP) data at 0.17 au. Using the Morlet wavelet transform, which is a good compromise between time and frequency resolution, we detect intermittent events that cover a wide range of scales. We show that these events correspond to embedded multi-scale structures, from MHD to sub-ion scales. Then, we study in more details the nature of these structures, which cover the whole turbulent cascade.
@@
The article is organized as follows: Section 2 describes the PSP data used in the analysis. In Section 3 we identify MHD, ion and sub-ion ranges of scales based on the magnetic spectral properties and positions of ion characteristic scales in the spectrum. Section 4 is dedicated to the detection method of coherent structures. Section 5 presents several theoretical models of the structures we think we cross by PSP at different scales. Then, we study the sensitivity of minimum variance results for different spacecraft trajectories across the model structures for different noise levels. In section 6 we describe a few examples of structures detected simultaneously at MHD, ion, and sub-ion scales. Section 7 describes the statistical study of the observed coherent events during 5h of the 1st PSP encounter at MHD, ion, and sub-ion scales. In section 8 we summarize and discuss the results.
@@ 2 Data
Figure 1: Overview of the solar wind data during the first perihelion of PSP at 0.17 au on November 6, 2018, between 00:00 and 05:00 UT. From top to bottom: (a) the magnetic field components in the RTN reference frame and the magnetic field modulus, (b) proton velocity radial component $V_{R}$, (c) angle between the magnetic field and the radial direction $\Theta_{\mathbf{BR}}$, (d) electron density $N_{e}$, (e) proton temperature $T_{p}$ (in blue) and core electron temperatures $T_{e,core}$ (in black), (f) core electron and proton plasma beta, (g) ion characteristic scales, $d_{i}$ and $\rho_{i}$.
We analyze a 5 hour time interval during the first perihelion, on November 6, 2018, [00:00, 05:00] UT, when the spacecraft at the distance of 0.17 au from the Sun measured the solar wind emerging from the small equatorial coronal hole (Bale et al., 2019; Kasper et al., 2019). The magnetic field during the chosen time interval is particularly highly-disturbed due to the presence of high-amplitude structures (including switchbacks, Bale et al., 2019; Perrone et al., 2020). The duration of the chosen interval is long enough to resolve the inertial range of MHD turbulence, but not too long, so that the PSP is magnetically connected to the same coronal hole and the PSP position is nearly at the same radial distance from the Sun.
We use the merged magnetic field measurements of two magnetometers: FIELDS/Fluxgate Magnetometer and Search Coil (Bale et al., 2016; Bowen et al., 2020). These data have a $3.4$ ms time resolution, which allows us to resolve a wide range of scales, from the MHD inertial range to the sub-ion range. Due to the anomaly of the Search Coil at one of its axes (which happened in March 2019), the full merged vector of the magnetic field is accessible only for the first perihelion (Bowen et al., 2020). Figure 1(a) shows the magnetic field magnitude $B(t)$ in black and three components in the RTN coordinate frame in color. The magnetic field vector fluctuates around $\langle\mathbf{B}\rangle=(-47,63,-5)$ nT. Its magnitude is nearly constant $|\mathbf{B}|=98\pm 5$ nT. The angle between the magnetic field and the radial direction changes from $25^{\circ}$ to $180^{\circ}$ as shown in Figure 1(c), with a dominance around quasi-perpendicular orientation, $\langle\Theta_{\mathbf{B}\mathbf{R}}\rangle=121^{\circ}$ with standard deviation $\sigma(\Theta_{\mathbf{B}\mathbf{R}})=26^{\circ}$.
To characterize plasma bulk velocity, we use the SWEAP/SPC Faraday cup instrument (Kasper et al., 2016). Proton velocity $\mathbf{V}$ is estimated from the 1st moment of the distribution function. Figure 1(b) shows its radial component $V_{R}$. The mean velocity is $\langle\mathbf{V}\rangle=(330,70,8)$ km/s, making a non-zero angle with the radial direction: $\langle\Theta_{\mathbf{VR}}\rangle=14^{\circ}$. This small deviation from the radial of the solar wind speed might not be accurate due to the field of view of the instrument together with the high angular satellite velocity during the considered time interval [Michael Stevens, private communication]. It may introduces a relative error of about 3% on the estimates of spatial scales $\ell$ using Taylor hypothesis.
@@
Proton temperature $T_{p}$ is estimated from the second moment of the distribution function measured by the SWEAP/SPC instrument. Figure 1(e) shows QTN electron core temperature $T_{e,core}$ (in black) and proton temperature $T_{p}$ (in blue). The mean temperatures are $\langle T_{e,core}\rangle=28.5$ eV and $\langle T_{p}\rangle=15$ eV.
We consider solar wind fluctuations over a wide range of timescales. The minimum time scale is determined by the temporal resolution of the magnetic data $dt=3.4\cdot 10^{-3}$ s, and the maximum by the total duration of the time interval $T=5$ h. The spatial scales of the corresponding range, estimated using the Taylor hypothesis with $V=340$ km/s, $\ell\in[1,10^{5}]$ km are much larger than the Debye length $\lambda_{D}\simeq 2$ m, then the plasma is quasi-neutral. During the analyzed time period, alpha particle abundance has not been measured, but it can be estimated as $A_{\alpha}=N_{\alpha}/N_{p}<5\%$ (Kasper et al., 2007; Alterman & Kasper, 2019; Liu et al., 2021). The quasi-thermal noise spectroscopy provides a more accurate measurement of the density than particle detectors, so we use $N_{p}=N_{e}$ and calculate proton plasma beta $\beta_{p}$ using the electron density: $\beta_{p}=N_{e}kT_{p}/(B^{2}/2\mu_{0})$, with $\mu_{0}$ being the magnetic permeability. Plasma beta for core electrons is defined as $\beta_{e,core}=N_{e}kT_{e,core}/(B^{2}/2\mu_{0})$. Both plasma $\beta$ parameters are well below unity as shown in Figure 1(f), $\langle\beta_{p}\rangle=0.22$, $\langle\beta_{e,core}\rangle=0.44$. Figure 1 (g) shows the time variation of the ion characteristic spatial scales $d_{i}=c/\omega_{pi}$, where $\omega_{pi}=\sqrt{4\pi N_{p}e^{2}/m_{p}}$ and $\rho_{i}=V_{th,\perp}/\omega_{ci}$, where $V_{th,\perp}=\sqrt{2k_{B}T_{\perp}/m_{p}}$ is the perpendicular proton thermal speed, $\omega_{ci}=eB/m_{p}c$ is the proton gyrofrequency.
3 MHD, ion, and sub-ion range identification
Figure 2: (a) Magnetic field total spectrum $S_{total}$ in orange and magnetic field modulus spectrum, that is the proxi of the parallel fluctuations spectrum $S_{\parallel}$ (Equation (5)), and (b) the ratio $S_{\parallel}/S_{total}$. The vertical lines show the characteristic ion scales: ion cyclotron frequency $f_{ci}$ (in blue), and the frequencies computed with the Doppler shifted ion gyroradius $f_{\rho i}$ (in black) and the Doppler shifted ion inertial length $f_{di}$ (in grey). The frequency ranges are highlighted: MHD in red, ion scales in green and sub-ion scales in blue.
First, we describe the magnetic field spectral properties of the analyzed time interval. We apply wavelet transform with Morlet mother function (Torrence & Compo, 1998):
@@ 3 MHD, ion, and sub-ion range identification
where $\omega_{0}=6$ is the angular frequency of oscillations in the mother function (with normalized time). The wavelet transform of the magnetic field component $B_{i}(t)$ is defined as the convolution of $B_{i}(t)$ with scaled, translated, and normalized $\psi_{0}(t)$ to have mother function $\psi$ with unit energy:
@@
where the sign indicates the complex conjugate.
Wavelet coefficients are influenced by the edge effects. The cone of influence (COI) curve separates the region of scales where edge effects become important as a function of time. To avoid this edge effect we consider a maximum scale equal to $\tau_{max}=10^{3}$ s. The intercept of $\tau_{max}$ with the COI curve determines the time sub-interval $T^{\prime}=$ [00:22:49, 04:37:11] UT, where wavelet coefficients at the scales $\tau<\tau_{max}$ are not influenced by the edge effect.
@@
where $\delta t=0.008$ s is the time-step of the PSP merged magnetic field data. The relation between Fourier frequencies $f$ and time scales $\tau$ is $f\simeq 1/\tau$ for the Morlet wavelets with $\omega_{0}=6$. In Figure 2(a), the blue line shows the PSD of compressive magnetic fluctuations. Compressive fluctuations are approximated here by the variation of magnetic field modulus. Indeed, this approximation is valid if the level of the fluctuations is significantly lower than the mean field $B_{0}$, i.e., $\delta B/B_{0}\ll 1$ (Perrone et al., 2016):
@@
In the inertial range and at higher frequencies the condition $\delta B/B_{0}\ll 1$ is valid. So we calculate the parallel PSD, $S_{\parallel}(\tau)$, as was done by Perrone et al. (2016):
@@
The ratio of compressible fluctuations to the total power spectral density $S_{\parallel}/S_{total}$ is shown in Figure 2(b). In the inertial range, parallel magnetic fluctuations are much less energetic than perpendicular ones ($\delta B_{\|}\ll\delta B_{\perp}$), as is usually observed in the solar wind. At the sub-ion scales, the fraction of the parallel $S_{\parallel}(\tau)/S_{total}(\tau)$ increases up to $\simeq 0.2$, which is consistent with the results of Salem et al. (2012) at 1 au. The authors suggested that the observed spectral ratio can be explained by the presence of the kinetic Alfvén wave (KAW) cascade with nearly perpendicular wavevectors ($k_{\perp}\gg k_{\|}$). However, analyzing Cluster measurements (Lacombe et al., 2017) and 2D hybrid numerical simulation (Matteini et al., 2020) found that the asymptotic compressibility value at sub-ion scales doesn’t match perfectly the KAW prediction. Finally, recent numerical simulations indicate that coherent structures, rather than waves, are energetically dominant on sub-ion scales (Papini et al., 2021).